I am very confused on this problem
The Danville Express travels 280 miles at a certain speed. If the speed had been increased by 5mph the trip could have been made in 1 hour less time find actual speed

Answer by scott8148(6628)

time = distance / rate

280 / x = [280 / (x + 5)] + 1

multiplying by x(x+5), the LCD ___ 280x + 1400 = 280x + x^2 + 5x

0 = x^2 + 5x - 1400

factoring ___ 0 = (x + 40)(x - 35)

x + 40 = 0 ___ x = -40 ___ negative value not realistic

x - 35 = 0 ___ x = 35